OpenStudy (anonymous):

List all real values of x such that f(x)=0. f(x)=((-17)/(x+18))+(-14/(x-9))

Parth (parthkohli):

Whatever the problem is(I don't understand). f(x) = blahabalahh Just equate it with zero. 0 = ((-17.........

OpenStudy (anonymous):

\[ -\frac{17}{x+18}-\frac{14}{x-9}=0 \]

OpenStudy (anonymous):

\[ \frac{-31 x-99}{(x-9) (x+18)}=0 \]

OpenStudy (campbell_st):

rewrite the equation \[\frac{17}{x + 18} = \frac{-14}{x - 9}\] cross multiply to eliminate the fractions \[17(x - 9) = -14(x + 18)\] 17x - 153 = -14x - 252 add 14 to both sides 31x - 153 = -252 add 153 to both sides 31x = -99 \[x = -\frac{99}{33}\]

OpenStudy (campbell_st):

oops \[x =- \frac{99}{31}\]

Parth (parthkohli):

Writing it like that was certainly a better way. Good job.

OpenStudy (anonymous):

thank you very much :)